Using an independent samples t-test a researcher obtains a t-statistic of 2.316, he rejects the null at the 2 tailed alpha level of 0.05.

Estimate N1+N2 (Total N, Total number of subjects):

less than 8

8 or more

10 or more

less than 10

7 or more
Answer #1

We are given the T statistic value to be 2.316 which is rejected at level 0.05. Since we do a two tailed test we consider the significance level to be 0.05/2= 0.025. Since the test is rejected so the critical value will be less than the T statistic value. From the T distribution table we see at df 8 that the value of t is less than 2.316. We know df = N1+N2-2. so the estimate of df = N1+N2= 10 . Fom the T table we see as the df increases from 8 the t value decreases than 2.316. So the correct estimate is third option , i.e, 10 or more.